# Orientability of Undirected Phylogenetic Networks to a Desired Class: Practical Algorithms and Application to Tree-Child Orientation

### v2

@@ 6 Experiments
We implemented our two proposed methods (Algorithms 1 and 2) and the existing exponential time algorithm described by Huber et al. (Algorithm 2 in [1]) using Python 3.11.6. In the implementation of Algorithms 1 and 2, we used the minimum_cycle_basis function from the Python package networkx to compute a minimal cycle basis. We note that the algorithm implemented in networkx is not the $O(|V||E|^{2}/\log|V|)$ algorithm given in [17] but the $O(|E|^{3}+|E||V|^{2}\log{|V|})$ algorithm given in Section 7.2 of [16]. Undirected binary phylogenetic networks were generated as test data using a method described in Appendix. The source code, test data, and the program used to generate the data are available at https://github.com/hayamizu-lab/tree-child-orienter. The full details of the results can be found at https://github.com/hayamizu-lab/tree-child-orienter/tree/main/results.
@@ 6.1 Experiment 1: Execution time
Figure 5: Results of Experiment 6.1. Comparison of the execution time of the three methods for tree-child orientable networks with $10$ leaves. Our exponential time algorithm (Algorithm 1) has been shown to be significantly faster than the existing exponential time algorithm described in [1].
@@ 6.3 Experiment 3: Accuracy for large graphs
In Experiment 6.3, we evaluated the performance of Algorithms 1 and 2 on the 471 larger sample networks with 20 leaves and 1 to 9 reticulations in Table 4. The experiment was performed on the same MacBook Pro used in Experiment 6.2. Since Algorithm 2 in [1] became infeasible for most cases with $r\geq 6$, we compared Algorithms 1 and 2 based on the number of tree-child orientations found and the time taken to find one.
@@
Table 1: Results of Experiment 6.2. The input networks have $10$ leaves. #YES (resp. #NO) is the number of tree-child orientable (resp. non-tree-child orientable) instances among the generated networks with each reticulation number $r$. #YES and #NO have been verified using the existing exponential time algorithm in [1]. The execution time here is the time taken to find a tree-child orientation or to output ‘NO’ or ‘Probably NO’.
@@
$r$YES Accuracy Execution Time (sec) Accuracy Execution Time (sec)
@@ 0 N/A N/A N/A N/A N/A N/A N/A N/A
Table 2: Results of Experiment 6.3. The input networks have $20$ leaves. #Graph is the number of generated networks with each reticulation number $r$. #YES is the number of YES instances among those networks, which is equal to the number of tree-child orientations found by Algorithm 1. #TC is the number of tree-child orientations found by Algorithm 2. The execution time here is the time taken to find a tree-child orientation.
@@
$r$Graph #YES Execution time (sec) #TC Execution Time (sec)
@@ 8 Conclusion and future work
The $C$-Orientation problem, which asks whether a given undirected binary phylogenetic network can be oriented to a directed phylogenetic network of a desired class $C$, is an important computational problem in phylogenetics. The complexity of this problem remains unknown for many network classes $C$, including the class of binary tree-child networks. A simple exponential time algorithm for $\mathcal{C}$-Orientation was provided in [1]. FPT algorithms for a special case of the problem were also proposed in [1], but their practical application is limited due to the intricate nature of the procedures and the challenges in implementation, despite the constraints imposed on $\mathcal{C}$. Additionally, no study has explored heuristic approaches to solve $\mathcal{C}$-Orientation in practice, even for a particular class $\mathcal{C}$ such as tree-child networks.
@@
Although we have used Tree-Child Orientation as a case study for performance evaluation, Algorithm 1 is applicable to orientation problems for classes $\mathcal{C}$ other than tree-child networks. For example, orientation for stack-free networks [22] or tree-based networks [23] is expected to be a good application, because one can quickly decide whether a given network belongs to such a class. Speeding up Algorithm 1 and extending it to non-binary networks are topics for future research. It would be also interesting to consider alternative FPT algorithms that do not use a cycle basis.
@@ 8 $\{13\}$
Supplementary information The source code, datasets and detailed experimental results are available at https://github.com/hayamizu-lab/tree-child-orienter (archived at swh:1:dir:666a10c01c14741702fddd5f8704b30bc90299e5).
@@
Acknowledgements We thank Yufeng Wu and Louxin Zhang for sharing code for randomly generating graphs, which we adapted for our experiments. We also appreciate the anonymous reviewers’ careful reading of our manuscript and useful comments.
[…]
Declarations
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Competing interests
The authors declare that they have no competing interests.
[…]
Funding MH is supported by JST FOREST Program Grant Number JPMJFR2135, Japan.

### v3

@@ 6 Experiments
We implemented our two proposed methods (the Tree-Child Orientation version of Algorithm 1, and Algorithm 2) and the existing exponential time algorithm described by Huber et al. (Algorithm 2 in [2]) using Python 3.11.6. In the implementation of Algorithms 1 and 2, we used the minimum_cycle_basis function from the Python package networkx to compute a minimal cycle basis. We note that the algorithm implemented in networkx is not the $O(|V||E|^{2}/\log|V|)$ algorithm given in [22] but the $O(|E|^{3}+|E||V|^{2}\log{|V|})$ algorithm given in Section 7.2 of [21]. Undirected binary phylogenetic networks were generated as test data using a method described in Appendix. The source code, test data, and the program used to generate the data are available at https://github.com/hayamizu-lab/tree-child-orienter. The full details of the results can be found at https://github.com/hayamizu-lab/tree-child-orienter/tree/main/results.
@@ 6.1 Experiment 1: Execution time
Figure 5: Results of Experiment 6.1. Comparison of the execution time of the three methods for tree-child orientable networks with 1010 leaves. Our exponential time algorithm (Algorithm 1) has been shown to be significantly faster than the existing exponential time algorithm described in [2].
@@ 6.3 Experiment 3: Accuracy for large graphs
In Experiment 6.3, we evaluated the performance of Algorithms 1 and 2 on the 471 larger sample networks with 20 leaves and 1 to 9 reticulations in Table 4. The experiment was performed on the same MacBook Pro used in Experiment 6.2. Since Algorithm 2 in [2] became intractable for most cases with $r\geq 6$, we compared Algorithms 1 and 2 based on the number of tree-child orientations found and the time taken to find one.
@@
Table 1: Results of Experiment 6.2. The input networks have 1010 leaves. #YES (resp. #NO) is the number of tree-child orientable (resp. non-tree-child orientable) instances among the generated networks with each reticulation number $r$. #YES and #NO have been verified using the existing exponential time algorithm in [2]. The execution time here is the time taken to find a tree-child orientation or to output ‘NO’ or ‘Probably NO’.
@@
$r$ #YES Accuracy Execution Time (sec) Accuracy Execution Time (sec)
@@ 0 N/A N/A N/A N/A N/A N/A N/A N/A
Table 2: Results of Experiment 6.3. The input networks have 2020 leaves. #Graph is the number of generated networks with each reticulation number $r$. #YES is the number of YES instances among those networks, which is equal to the number of tree-child orientations found by Algorithm 1. #TC is the number of tree-child orientations found by Algorithm 2. The execution time here is the time taken to find a tree-child orientation.
@@
$r$ #Graph #YES Execution time (sec) #TC Execution Time (sec)
@@ 7 Discussion
As we have seen in Section 6, the heuristic method (Algorithm 2) is significantly faster than the exact FPT method (Algorithm 1), but it suffers from an increasing rate of false negatives as the reticulation number $r$ grows. We therefore examine the theoretical reasons behind its limitations and explore its potential biological applications.
@@ 8 Conclusion and future work
The $\mathcal{C}$-Orientation problem, which asks whether a given undirected binary phylogenetic network can be oriented to a directed phylogenetic network of a desired class $\mathcal{C}$, is an important computational problem in phylogenetics. The complexity of this problem remains unknown for many network classes $\mathcal{C}$, including the class of binary tree-child networks [4, 5]. A simple exponential time algorithm for $\mathcal{C}$-Orientation was provided in [2]. FPT algorithms for a special case of the problem were also proposed in [2], but their practical application is limited due to the intricate nature of the procedures and the challenges in implementation, despite the constraints imposed on $\mathcal{C}$. Additionally, no study has explored heuristic approaches to solve $\mathcal{C}$-Orientation in practice, even for a particular class $\mathcal{C}$ such as tree-child networks.
@@
Although we have used Tree-Child Orientation as a case study for performance evaluation, Algorithm 1 is applicable to orientation problems for any classes $\mathcal{C}$ other than tree-child networks as long as the membership test is tractable. For example, orientation for stack-free networks [27] or tree-based networks [28] is expected to be a good application, because one can quickly decide whether a given network belongs to such a class. Speeding up Algorithm 1 and extending it to non-binary networks are topics for future research. It would be also interesting to consider alternative FPT algorithms that do not use a cycle basis.